# A cross-platform analysis of polarization and echo chambers in climate change discussions

### v2

@@ Appendix C Additional results polarization and entropy in Twitter
In Fig. S10, we show the $20$ most tweeted and retweeted hashtags in the two communities analyzed per network. We observe a large overlap in the most tweeted hashtags between communities with a dominance of the generalistic ones.
Figure S11: Hashtag similarity in the COP27, IPCC, Climate Refugees and Doñana datasets. Similarity matrix between the set of top 20 hashtags in each of the two opposed communities. Results for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
Finally, we employ the same approach used to study the chamber overlap to characterize and compare the use of hashtags (Fig. S11). We calculated the weighted Jaccard similarity between hashtags on the fraction of tweets by community. We observe a clear grouping of the hashtags, which is much stronger in the Doñana network, indicating a stronger polarization in the use of hashtags.
Appendix D Additional results Youtube
We show in Fig. S12 the network between YouTube posts for each of the datasets analyzed.
@@ Appendix D Additional results Youtube
Figure S12: Network of Youtube posts extracted from the Twitter datasets. Each node corresponds to a post and links to the number of co-commenting users. Each node is colored according to the community calculated by optimizing the modularity [75].
We have analyzed the subnetwork focusing on those communities spreading skeptical content on climate change and the larger community with mainstream content, and we report them in Fig. S13.
@@
Figure S13: Network of YouTube posts after selecting the two main communities of mainstream and skeptics. Interaction between the posts after selecting the larger communities with mainstream and skeptical ideas. We calculated the communities by optimizing the modularity for two partitions. We identified the positioning of the communities by inspecting the videos.
Figure S14: Network of YouTube users. Networks of interactions between YouTube users. The weight of links is the number of common posts, and the color corresponds to the community calculated by the optimization of modularity [75].
Figure S15: Polarization in the network of Youtube users. Quantification of the modularity and EI indices in the user network. We have calculated the indices in the real network and in 100 randomizations where we preserve the degree sequences.

### v3

@@ Appendix C Policy and affective polarisation
Figure S16: Hashtag proportion test. We have conducted the ztest between the proportion of hashtags in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks. Blue bars indicate significance (p-value<<0.05).
S3 Topic and sentiment analysis
We have implemented a Latend Dirichlet Allocation model (LDA) to uncover the main topics discussed in each of the datasets. In Fig. S17 we report the most important words for each of the topics detected in the datasets. We have analyzed the polarisation around each of the topics Fig. S18.
Figure S17: Topics detected in the networks analized. Results of the Latent Dirichlet Allocation model (LDA) for eight topics in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks.
Figure S18: Distribution of the polarisation around the discussion topics. Topics detected and polarisation in (a) COP27, (b) IPCC,(c) Climate Refugees, (d) Doñana. Blue bars indicate that the frequency of topics between communities is statistically significant.
We have implemented a sentiment analysis on the four Twitter datasets using a Roberta-based model fine-tuned on around 58 million of Tweets [77]. We provide the Mann-Whitney U tests on the quoted messages in Fig. S19.
Figure S19: Mann-Whitney U test in quote sentiment. We have conducted the Mann-Whitney U test of the negative sentiment between communities in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks. The test was conducted to assess if the entries in the vertical axes are greater than those in the horizontal axes.
Appendix D Cross-platform analysis
We present the network of YouTube posts extracted from the Twitter datasets, where each node represents a post and is linked based on the number of shared commenters Figure S20. Nodes are color-coded according to community structures identified through modularity optimization [69]. To further investigate the network’s structure, we analyze a subnetwork focusing on communities that predominantly spread skeptical narratives about climate change, as well as the largest community associated with mainstream content. The results of this analysis are depicted in Fig. S21.
@@ Appendix D Cross-platform analysis
Figure S20: Network of Youtube posts extracted from the Twitter datasets. Each node corresponds to a post and links to the number of co-commenting users. Each node is coloured according to the community calculated by optimising the modularity [69].
@@
Figure S21: Network of YouTube posts after selecting the two main communities of mainstream and low-reliability content. Interaction between the posts after selecting the larger communities with mainstream and low-reliability content. We calculated the communities by optimizing the modularity for two partitions. We identified the positioning of the communities by inspecting the videos.
In Fig S22 we show the Twitter subnetwork of users that shared YouTube links related to climate change coloured according to the YouTube community of the videos. Only connected components with more than 5 users are depicted.
Figure S22: Network between Twitter users coloured by YouTube community. Interactions between Twitter users that have published YouTube Content in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. Each dot is coloured by the content of YouTube communities they have tweeted on.
We have computed the probability $P(C_{i},C_{j})$ of finding a user that posted a video in YouTube community $C_{i}$ connected to a user that posted a video in YouTube community $C_{j}$. To account for network effects, we have calculated $\widetilde{P}(C_{i},C_{j})=P(C_{i},C_{j})/P^{\rm null}(C_{i},C_{j})$ normalizing the observed probabilities by the corresponding null values over 100 random reshufflings of YouTube communities (Fig. S23)
Figure S23: YouTube Community correlations in the Twitter network. Probability of finding a user than posted a video in a YouTube community $C_{i}$ next to a user that posted a video in a YouTube community $C_{j}$. The results are normalized by 1000 random reshufflings of YouTube communities. The results correspond to (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks.
network community channel name viewCount subscriberCount videoCount
cop27 0 ABC News 14083290454 15900000 86239
cop27 0 TEDx Talks 7712122864 40100000 210266
cop27 0 WION 4701543170 8470000 136298
cop27 0 Al Jazeera English 4158638169 12800000 111285
cop27 1 ABC News 14083290454 15900000 86239
cop27 1 BBC News Hindi 7893209075 17800000 24225
cop27 1 BBC News 5099185254 15600000 20649
cop27 1 Hindustan Times 4842818648 6530000 58180
cop27 2 WION 4701543170 8470000 136298
cop27 2 Sky News Australia 4140464784 3900000 137500
cop27 2 CNN-News18 2967286345 6660000 160062
cop27 2 euronews 1368026874 2320000 99119
cop27 3 Democracy Now! 672210420 1980000 15630
cop27 3 BreakThrough News 138428175 609000 1403
cop27 3 CNBC International TV 116672225 362000 23849
cop27 3 Earthling Ed 54396530 452000 333
cop27 4 UOL 3147321219 3880000 62664
cop27 4 CNN Brasil 2660548853 4780000 79898
cop27 4 Poder360 441840465 1300000 19396
cop27 4 United Nations 406648137 2870000 16539
cop27 5 DW Español 2280313778 4150000 39964
cop27 5 EL PAÍS 1525037481 2550000 32904
cop27 5 RTVE Noticias 1378481206 2020000 25701
cop27 5 Presidencia de la República - Colombia 92079186 318000 20699
donana 0 Negocios TV 961486669 1010000 51192
donana 0 COPE 428257253 680000 6994
donana 0 LibertadDigital 332062887 684000 14666
donana 0 DISTRITOTV 172849322 331000 17512
donana 1 Canal Red 40410031 309000 700
donana 1 Pandemia Digital 21969182 77300 1701
donana 2 EL PAÍS 1525037481 2550000 32904
donana 2 RTVE Noticias 1378481206 2020000 25701
donana 2 elDiarioes 172800438 320000 9763
donana 2 Hoy por Hoy 37690183 71900 2893
ipcc 0 ABC News 14083290454 15900000 86239
ipcc 0 MSNBC 12050180260 6440000 77964
ipcc 0 BBC News 5099185254 15600000 20649
ipcc 0 Vox 3427290494 11900000 1677
ipcc 1 Fox News 16126544384 11000000 101812
ipcc 1 NBC News 6952022416 9650000 61006
ipcc 1 Sky News Australia 4140464784 3900000 137500
ipcc 1 PragerU 1898459165 3210000 3188
ipcc 2 Sky News Australia 4140464784 3900000 137500
ipcc 2 Sky News 4139448980 7370000 44104
ipcc 2 GBNews 1199821224 1170000 58097
ipcc 2 Rebel News 716906956 1640000 21112
ipcc 3 Fox News 16126544384 11000000 101812
ipcc 3 The Hill 1462188609 1900000 30762
ipcc 3 KTLA 5 1023133916 1020000 22849
ipcc 3 WUSA9 337193118 432000 52569
ipcc 4 WDR aktuell 83165245 103000 1139
ipcc 4 Carsten Jahn - TEAM HEIMAT 76120963 179000 879
ipcc 4 Stuttgarter Zeitung 60081195 60000 3424
ipcc 4 eingeSCHENKt.tv 19415430 131000 544
ipcc 5 De Nieuwe Wereld 26340396 101000 1797
ipcc 5 1000frolly PhD 10920557 30100 170
ipcc 5 Holland Gold 2880125 19800 181
ipcc 5 Tom Nelson 1587268 12400 219
refugees 0 Fox News 16126544384 11000000 101812
refugees 0 MSNBC 12050180260 6440000 77964
refugees 0 TEDx Talks 7712122864 40100000 210266
refugees 0 NBC News 6952022416 9650000 61006
refugees 1 CNN 15327838440 16000000 163405
refugees 1 ABC News 14083290454 15900000 86239
refugees 1 NBC News 6952022416 9650000 61006
refugees 1 BBC News 5099185254 15600000 20649
refugees 2 Fox News 16126544384 11000000 101812
refugees 2 MSNBC 12050180260 6440000 77964
refugees 2 NBC News 6952022416 9650000 61006
refugees 2 Sky News Australia 4140464784 3900000 137500
refugees 3 AllatRa TV English 2840110 29500 964
refugees 3 Creative Society 2664803 40600 1163
refugees 3 ALLATRA TV International 1507255 18500 982
refugees 4 VICE News 3156884335 8830000 6480
refugees 4 Global News 2490061489 4110000 37952
refugees 4 DW News 2241829454 5010000 33064
refugees 4 CBC News 1955874636 3450000 28005
refugees 5 TEDx Talks 7712122864 40100000 210266
refugees 5 NBC News 6952022416 9650000 61006
refugees 5 Democracy Now! 672210420 1980000 15630
Table S11: Top 4 most relevant YouTube channels by view count in the 5 largest communities.